OpenStudy (anonymous):

Use the distributive property to rewrite the expression and solve. 9 72 = 9 + 9 2 =

OpenStudy (anonymous):

help please

OpenStudy (anonymous):

Sorry...I don't get what you were trying to write...could you try to write it using the equation tool?

OpenStudy (anonymous):

Use the distributive property to rewrite the expression and solve. 9 * 72 = * + 9 + 9 * 2 =

OpenStudy (fibonaccichick666):

Distributive property implies some parentheses somewhere

OpenStudy (fibonaccichick666):

ok so you have \[9(72)=x*y+9+9*2\]

OpenStudy (anonymous):

yes

OpenStudy (fibonaccichick666):

so let's start by breaking down 72 by itself

OpenStudy (anonymous):

ok

OpenStudy (fibonaccichick666):

so, how else can you write 72?

OpenStudy (anonymous):

7.2

OpenStudy (fibonaccichick666):

72=7.2?

OpenStudy (anonymous):

i dont get it

OpenStudy (fibonaccichick666):

Try thinking about 72 as a sum, find 3 numbers you can add together to get 72

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

23 24 25

OpenStudy (anonymous):

@FibonacciChick666

OpenStudy (anonymous):

@HailKK

OpenStudy (anonymous):

@Holly00d1248

OpenStudy (fibonaccichick666):

ok, so that works, now, let's apply that concept and think about how the distributive property works, so if you have \[6(x+y)=?\] So can you distribute the 6?

OpenStudy (anonymous):

so what plus what = 6

OpenStudy (fibonaccichick666):

no, I'm sorry, I do not understand that statement

OpenStudy (anonymous):

what do i do

OpenStudy (fibonaccichick666):

please read about the distributive property here http://www.purplemath.com/modules/simparen.htm

OpenStudy (anonymous):

@FateRocks8D

OpenStudy (anonymous):

um , sorry love im not good with distributive property I failed it a few time sbut I think @Nnesha could see about this

OpenStudy (anonymous):

@Nnesha

OpenStudy (anonymous):

@Destinymasha

OpenStudy (anonymous):

@TheOcean

OpenStudy (anonymous):

@RaeBeach9296

OpenStudy (anonymous):

help please medal and fan

OpenStudy (anonymous):

was up

OpenStudy (anonymous):

@jdoe0001 @undeadknight26 @Maria195 hope u guys know this o_o

OpenStudy (anonymous):

u need t find the *?

OpenStudy (anonymous):

what is in the blanks

OpenStudy (anonymous):

9 * 72 = * + 9 + 9 * 2 = hmm... i know how to solve one so solve one first then the second one

OpenStudy (anonymous):

ok

OpenStudy (anonymous):